Find the value of x and KL if K is between J and L. JK=2x, KL=x+2, and JL=5x-10.
Thank u :)

Answer by Cromlix(4381)

Hi there,
If K lies between J and L
then JK + KL = JL
2x + (x + 2) = 5x - 10
2x + x + 2 = 5x - 10
Collect like terms:
2x + x - 5x = -10 - 2
-2x = -12
x = 6
Hope this helps:-)